OpenStudy (anonymous):

4x^2+x−3=0, solve by factoring. The solutions are 3/4 or -1 please help solve?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

compare the eqn with ax^2+bx+c=0 then we know 

OpenStudy (anonymous):

can u understand now?

OpenStudy (anonymous):

i understand what yo're saying, but in my book it shows that thesolutions are 3/4 or -1, but i just dont understand how they got that?

OpenStudy (anonymous):

you're*

OpenStudy (anonymous):

now i think u hv no any problem on such type of problems

OpenStudy (anonymous):

4x^2+x−3=(4x-3)(x+1)=0,the solutions are x=3/4 or x=-1.You are right!

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

Mathematical cross method

OpenStudy (anonymous):

do you know the box method... it is where you take the factors of the first term, and the factors of the last term, and when you cross multiply those factors you get a product of the middle term.

OpenStudy (anonymous):